# SuperM2M: Supervised and Mixture-to-Mixture Co-Learning for Speech Enhancement and Robust ASR

### v2

@@ 4.5 Weaknesses of M2M Training
5 SuperM2M
The previous section points out that M2M suffers from source and frequency permutation, and source ambiguity. On the other hand, although M2M training can be performed on real mixtures, there may not be many paired close-talk and far-field real-recorded mixtures available, as collecting real data is effort-consuming. In comparison, supervised models can be readily trained on massive simulated mixtures, as one can easily simulate as many mixtures as one considers sufficient. In addition, they do not suffer from source and frequency permutation, and source ambiguity.
@@ 5.1 Supervised and Mixture-to-Mixture Co-Learning
In detail, at each training step, we sample either a mini-batch of real close-talk and far-field mixture pairs or a mini-batch of simulated far-field mixtures for DNN training. The loss on real data is $\mathcal{L}_{\text{MC}}$ in (5), and the loss on simulated data is
@@
$\mathcal{L}_{\text{Target},q}$ $=\frac{1}{\sum\nolimits_{t,f}|Y_{q}(t,f)|}\sum_{t,f}\mathcal{G}\Big{(}S_{q}(t,f),\hat{S}_{q}(t,f)\Big{)},$ (15)
@@
where $\mathcal{G}(\cdot,\cdot)$ is defined in (10), $S_{q}$ and $V_{q}$ are obtained through simulation, and the denominator balances the loss values with the ones in M2M training.
Figure 3: Robust ASR pipeline, where enhanced speech $\hat{s}_{q}=\text{iSTFT}(\hat{S}_{q})$ is fed to backend ASR models for recognition. No joint training is performed. An optional speaker reinforcement module [72], which adds a scaled version of the input mixture signal $y_{q}$ to $\hat{s}_{q}$, can be included.
@@ 6 Experimental Setup
We do not use $\hat{S}_{q}$ to derive linear beamforming results for ASR [1, 33, 43], although this has been extremely popular, as we would like to validate whether the enhcanced speech $\hat{S}_{q}$ itself is close to target speech and whether $\hat{S}_{q}$ itself can yield better ASR performance. We do not jointly train enhancement models with ASR models, as this requires knowledge of ASR models and would not accurately reflect the accuracy of $\hat{S}_{q}$ itself. We aim at building enhancement models that can produce enhanced speech with low distortion and high reduction to non-target signals. This way, the enhancement models could improve the robustness of many subsequent applications, not just limited to ASR.
In a nutshell, our main goal is to show, through SuperM2M, whether $\hat{S}_{q}$ itself would be better on real test data. We validate SuperM2M on CHiME-4 [42], a dataset consisting of simulated mixtures and real-recorded close-talk and far-field mixture pairs. To further show the effectiveness and potential of SuperM2M, a minor goal is to show whether SuperM2M can lead to state-of-the-art ASR performance on CHiME-4.
The rest of this section describes the CHiME-4 dataset, miscellaneous system configurations, comparison systems, evaluation metrics, and several tricks to improve robust ASR performance.
@@ 6.1 CHiME-4 Dataset
CHiME-4 [42] is a major corpus for evaluating robust ASR and speech enhancement algorithms. It is recorded by using a tablet mounted with $6$ microphones, with the second microphone on the rear and the others facing front. The signals are recorded in four representative environments (including cafeteria, buses, pedestrian areas, and streets), where reverberation and directional, diffuse, transient and non-stationary noises naturally exist. During data collection, the target speaker hand-holds the tablet in a designated environment, and reads text prompts shown on the screen of the tablet. The target speaker wears a close-talk microphone so that the close-talk mixture can be recorded at the same time along with far-field mixtures recorded by the microphones on the tablet. The number of simulated and real-recorded utterances is listed in Table 2.
@@ 6.2 Evaluation Setup - Robust ASR
We check whether SuperM2M can improve ASR performance by feeding its enhanced speech to ASR models for decoding, following the pipeline in Fig. 3. We consider two ASR models. The first one is Whisper Large v2444https://huggingface.co/openai/whisper-large-v2 [73], pre-trained on massive data. We use its text normalizer to normalize hypothesis and reference text before computing WER. The second one is trained on the official CHiME-4 mixtures plus the clean signals in WSJ00 by using the public recipe [53] [^5] https://github.com/espnet/espnet/blob/master/egs2/chime4/asr1/conf/tuning/train_asr_transformer_wavlm_lr1e-3_specaug_accum1_preenc128_warmup20k.yaml in ESPnet. It is an encoder-decoder transformer-based model, trained on WavLM features [74] and using a transformer language model in decoding. Note that the WERs computed by ESPnet should not be directly compared with the ones by Whisper due to different text normalization.
@@ 6.5 Miscellaneous Configurations
TF-GridNet [18], which has shown strong separation performance in major benchmarks in supervised speech separation, is used as the DNN architecture. We consider two setups. Using the symbols defined in Table I of [18], the first one (denoted as TFGridNetv1) sets its hyper-parameters to $D=100$, $B=4$, $I=2$, $J=2$, $H=200$, $L=4$ and $E=2$, and the second one (denoted as TFGridNetv2) to $D=128$, $B=4$, $I=1$, $J=1$, $H=200$, $L=4$ and $E=4$. Please do not confuse these symbols with the ones in this paper. The models have $\sim$6.36.36.36.3 and $\sim$5.45.45.45.4 million parameters respectively. The v1 model uses around half of the computation and memory of v2, and is utilized for faster experimentation.
@@ 6.6 Comparison Systems
We consider the same DNN model trained only on the simulated data of CHiME-4 via supervised learning as the major baseline for comparison. We use exactly the same configurations as that in SuperM2M for traning. We denote this baseline as Supervised, to differentiate it with SuperM2M. Since CHiME-4 is a public and pupular dataset, many existing models can be used directly for comparison.
@@ 6.7.1 SNR Augmentation for Simulated Training Mixtures
6.7.2 iSTFT-STFT Projection
We apply inverse STFT (iSTFT) followed by STFT operations to the DNN estimates $\hat{S}_{q}$ and $\hat{V}_{q}$ before loss computation, i.e., $\hat{S}_{q}\coloneqq\text{STFT}(\text{iSTFT}(\hat{S}_{q}))$ and $\hat{V}_{q}\coloneqq\text{STFT}(\text{iSTFT}(\hat{V}_{q}))$. See Fig. 2 for an illustration. This often yields slight improvement, as the losses now penalize the RI components and magnitudes extracted from re-synthesized signals, which are the final system output used for human hearing and downstream tasks [79, 80, 69]. Notice that, in Fig. 3, ASR features are extracted from re-synthesized signals.
@@
(#input mics: $1$; ASR model: Whisper Large v2)
[…]
2 Supervised S iNeuBe [13] - - - - - $15.1$ $0.954$
4d SuperM2M S+R TFGridNetv1 5 6+1 - 16.901 048 553 260 8916.9010485532608916.901\,048\,553\,260\,8916.901 048 553 260 89 17.357 574 660 580 04517.35757466058004517.357\,574\,660\,580\,04517.357 574 660 580 045 2.362 516 766 335 0252.3625167663350252.362\,516\,766\,335\,0252.362 516 766 335 025 0.959 485 302 685 890 20.95948530268589020.959\,485\,302\,685\,890\,20.959 485 302 685 890 2 7.341 077 767 021 627.341077767021627.341\,077\,767\,021\,627.341 077 767 021 62 5.288 209 430 841 8385.2882094308418385.288\,209\,430\,841\,8385.288 209 430 841 838 12.226 013 234 077 7512.2260132340777512.226\,013\,234\,077\,7512.226 013 234 077 75 7.507 367 767 954 0877.5073677679540877.507\,367\,767\,954\,0877.507 367 767 954 087
4e SuperM2M S+R TFGridNetv1 6 6+1 - 16.859 340 666 821 5116.8593406668215116.859\,340\,666\,821\,5116.859 340 666 821 51 17.379 205 978 436 4717.3792059784364717.379\,205\,978\,436\,4717.379 205 978 436 47 2.414 184 342 550 5092.4141843425505092.414\,184\,342\,550\,5092.414 184 342 550 509 0.961 652 581 435 139 40.96165258143513940.961\,652\,581\,435\,139\,40.961 652 581 435 139 4 7.341 077 767 021 627.341077767021627.341\,077\,767\,021\,627.341 077 767 021 62 5.288 209 430 841 8385.2882094308418385.288\,209\,430\,841\,8385.288 209 430 841 838 12.226 013 234 077 7512.2260132340777512.226\,013\,234\,077\,7512.226 013 234 077 75 7.507 367 767 954 0877.5073677679540877.507\,367\,767\,954\,0877.507 367 767 954 087
4f SuperM2M S+R TFGridNetv1 7 6+1 - 16.963 124 431 443 93716.96312443144393716.963\,124\,431\,443\,93716.963 124 431 443 937 17.398 348 091 374 1717.3983480913741717.398\,348\,091\,374\,1717.398 348 091 374 17 2.420 927 345 301 165 72.42092734530116572.420\,927\,345\,301\,165\,72.420 927 345 301 165 7 0.961 293 461 207 607 60.96129346120760760.961\,293\,461\,207\,607\,60.961 293 461 207 607 6 7.228 138 109 067 4417.2281381090674417.228\,138\,109\,067\,4417.228 138 109 067 441 5.021 983 784 437 9015.0219837844379015.021\,983\,784\,437\,9015.021 983 784 437 901 12.194 995 864 350 70312.19499586435070312.194\,995\,864\,350\,70312.194 995 864 350 703 6.902 435 241 197 4566.9024352411974566.902\,435\,241\,197\,4566.902 435 241 197 456
4g SuperM2M S+R TFGridNetv1 8 6+1 - 16.627 877 601 619 9216.6278776016199216.627\,877\,601\,619\,9216.627 877 601 619 92 17.397 258 721 839 1217.3972587218391217.397\,258\,721\,839\,1217.397 258 721 839 12 2.414 387 100 934 9822.4143871009349822.414\,387\,100\,934\,9822.414 387 100 934 982 0.960 830 590 588 033 90.96083059058803390.960\,830\,590\,588\,033\,90.960 830 590 588 033 9 7.260 406 582 768 6357.2604065827686357.260\,406\,582\,768\,6357.260 406 582 768 635 5.026 017 506 353 1125.0260175063531125.026\,017\,506\,353\,1125.026 017 506 353 112 12.045 078 577 336 64112.04507857733664112.045\,078\,577\,336\,64112.045 078 577 336 641 6.954 138 875 962 986.954138875962986.954\,138\,875\,962\,986.954 138 875 962 98
[…]
7 SuperM2M S+R TFGridNetv2 est. 6+1 - 16.605 623 853 251 7916.6056238532517916.605\,623\,853\,251\,7916.605 623 853 251 79 17.418 666 160 478 68317.41866616047868317.418\,666\,160\,478\,68317.418 666 160 478 683 2.510 620 386 221 1922.5106203862211922.510\,620\,386\,221\,1922.510 620 386 221 192 0.962 660 024 574 209 50.96266002457420950.962\,660\,024\,574\,209\,50.962 660 024 574 209 5 7.046 627 944 498 2257.0466279444982257.046\,627\,944\,498\,2257.046 627 944 498 225 4.783 994 191 440 4424.7839941914404424.783\,994\,191\,440\,4424.783 994 191 440 442 11.342 018 196 856 90711.34201819685690711.342\,018\,196\,856\,90711.342 018 196 856 907 6.023 473 450 183 5486.0234734501835486.023\,473\,450\,183\,5486.023 473 450 183 548
[…]
9
Close-talk
Mixture
6.7.3 Run-Time Speaker Reinforcement for Robust ASR
At run time, in default we feed $\hat{s}_{q}$ for ASR. Alternatively, we employ a technique named speaker reinforcement [72], where $\hat{s}_{q}$ is re-mixed with the input mixture $y_{q}$ at an energy level of $\gamma$ dB before recognition. See Fig. 3 for an illustration. That is, $\hat{s}_{q}+\eta\times y_{q}$, where $\eta\in{\mathbb{R}}_{>0}$ and $\gamma=10\times\text{log}_{10}(\|\hat{s}_{q}\|^{2}/\|\eta\times y_{q}\|^{2})$. We find this technique usually effective for ASR, as the re-mixed input mixture can alleviate distortion to target speech.
7 Evaluation Results
This section reports our evaluations results on CHiME-4. Following earlier studies [35, 40], we set $\dot{I}$ and $\ddot{I}$ (described in Table 1) to $20$, and $\dot{J}$ to $1$. $\ddot{J}$ can be tuned to a value shared for all the training mixtures, or it can be estimated, on the fly, for each training mixture by using the method described in Section 4.4. Since CHiME-4 is known to contain minor reverberation, we set both $K$ and $\Delta$ to 00, meaning that $\widetilde{\hat{\mathbf{S}}}_{q}(t,f)$ and $\hat{\mathbf{g}}_{q}(f)$ in (6) do not exist. All the six far-field microphones are used for computing $\mathcal{L}_{\text{MC}}$ in (5). We emphasize that we spent minimal amount of effort on hyper-parameter tuning. Other hyper-parameter setup could lead to better performance.
Table 4: SuperM2M vs. purely-supervised models on CHiME-4 (#input mics: $6$; ASR model: Whisper Large v2)
@@ 7 Evaluation Results
4d SuperM2M S+R TFGridNetv1 5 6+1 - 22.278 870 459 758 87422.27887045975887422.278\,870\,459\,758\,87422.278 870 459 758 874 22.600 104 856 813 6122.6001048568136122.600\,104\,856\,813\,6122.600 104 856 813 61 3.070 647 907 076 463.070647907076463.070\,647\,907\,076\,463.070 647 907 076 46 0.984 785 720 375 6820.9847857203756820.984\,785\,720\,375\,6820.984 785 720 375 682 3.517 263 633 430 138 43.51726363343013843.517\,263\,633\,430\,138\,43.517 263 633 430 138 4 3.965 148 642 652 575 33.96514864265257533.965\,148\,642\,652\,575\,33.965 148 642 652 575 3 4.052 936 311 000 8284.0529363110008284.052\,936\,311\,000\,8284.052 936 311 000 828 4.156 972 235 148 1314.1569722351481314.156\,972\,235\,148\,1314.156 972 235 148 131
4e SuperM2M S+R TFGridNetv1 6 6+1 - 22.339 721 592 086 73222.33972159208673222.339\,721\,592\,086\,73222.339 721 592 086 732 22.692 359 542 666 7922.6923595426667922.692\,359\,542\,666\,7922.692 359 542 666 79 3.101 081 421 429 4613.1010814214294613.101\,081\,421\,429\,4613.101 081 421 429 461 0.984 796 130 298 721 50.98479613029872150.984\,796\,130\,298\,721\,50.984 796 130 298 721 5 3.606 001 936 108 422 33.60600193610842233.606\,001\,936\,108\,422\,33.606 001 936 108 422 3 3.944 980 033 076 523.944980033076523.944\,980\,033\,076\,523.944 980 033 076 52 3.975 392 886 683 209 33.97539288668320933.975\,392\,886\,683\,209\,33.975 392 886 683 209 3 4.213 846 233 390 2084.2138462333902084.213\,846\,233\,390\,2084.213 846 233 390 208
4f SuperM2M S+R TFGridNetv1 7 6+1 - 22.205 550 543 286 58422.20555054328658422.205\,550\,543\,286\,58422.205 550 543 286 584 22.523 039 187 784 7922.5230391877847922.523\,039\,187\,784\,7922.523 039 187 784 79 3.096 138 221 567 327 43.09613822156732743.096\,138\,221\,567\,327\,43.096 138 221 567 327 4 0.984 329 144 290 849 90.98432914429084990.984\,329\,144\,290\,849\,90.984 329 144 290 849 9 3.529 364 311 068 086 63.52936431106808663.529\,364\,311\,068\,086\,63.529 364 311 068 086 6 3.993 384 696 059 053 83.99338469605905383.993\,384\,696\,059\,053\,83.993 384 696 059 053 8 4.006 410 256 410 2564.0064102564102564.006\,410\,256\,410\,2564.006 410 256 410 256 4.244 868 414 249 5224.2448684142495224.244\,868\,414\,249\,5224.244 868 414 249 522
4g SuperM2M S+R TFGridNetv1 8 6+1 - 22.356 131 915 793 27422.35613191579327422.356\,131\,915\,793\,27422.356 131 915 793 274 22.651 026 868 914 45722.65102686891445722.651\,026\,868\,914\,45722.651 026 868 914 457 3.135 788 249 608 242 33.13578824960824233.135\,788\,249\,608\,242\,33.135 788 249 608 242 3 0.984 945 777 711 785 30.98494577771178530.984\,945\,777\,711\,785\,30.984 945 777 711 785 3 3.509 196 515 004 843.509196515004843.509\,196\,515\,004\,843.509 196 515 004 84 3.928 845 145 415 6753.9288451454156753.928\,845\,145\,415\,6753.928 845 145 415 675 3.965 053 763 440 863.965053763440863.965\,053\,763\,440\,863.965 053 763 440 86 4.193 164 779 483 9984.1931647794839984.193\,164\,779\,483\,9984.193 164 779 483 998
@@
7 SuperM2M S+R TFGridNetv2 est. 6+1 - 22.849 200 136 282 22822.84920013628222822.849\,200\,136\,282\,22822.849 200 136 282 228 23.214 971 050 443 76523.21497105044376523.214\,971\,050\,443\,76523.214 971 050 443 765 3.215 787 200 223 3163.2157872002233163.215\,787\,200\,223\,3163.215 787 200 223 316 0.986 786 352 625 754 60.98678635262575460.986\,786\,352\,625\,754\,60.986 786 352 625 754 6 3.384 156 179 412 7143.3841561794127143.384\,156\,179\,412\,7143.384 156 179 412 714 3.807 833 487 959 340 53.80783348795934053.807\,833\,487\,959\,340\,53.807 833 487 959 340 5 3.773 779 983 457 4033.7737799834574033.773\,779\,983\,457\,4033.773 779 983 457 403 4.043 224 238 663 9784.0432242386639784.043\,224\,238\,663\,9784.043 224 238 663 978

### v3

@@ 6.2 Supervised and Mixture-to-Mixture Co-Learning
The previous subsection points out that M2M suffers from source and frequency permutation, and source ambiguity. On the other hand, although M2M training can be performed on real mixtures, there may not be many paired close-talk and far-field real-recorded mixtures available, as collecting real data is effort-consuming. In comparison, supervised models can be readily trained on massive simulated mixtures, as one can easily simulate as many mixtures as one considers sufficient. In addition, they do not suffer from source and frequency permutation, and source ambiguity.
@@
In detail, at each training step, we sample either a mini-batch of real close-talk and far-field mixture pairs or a mini-batch of simulated far-field mixtures for DNN training. The loss for the mini-batch of real data is $\mathcal{L}_{\text{MC}}$ in (4), and the loss for the mini-batch of simulated data is
@@
$\mathcal{L}_{\text{Target},q}$ $=\frac{1}{\sum\nolimits_{t,f}|Y_{q}(t,f)|}\sum_{t,f}\mathcal{G}\Big{(}X_{q}(t,f),\hat{X}_{q}(t,f)\Big{)},$ (13)
@@
where $\mathcal{G}(\cdot,\cdot)$ is defined in (8), $X_{q}$ and $V_{q}$ are obtained through simulation, and the denominator balances the loss values with the ones in M2M training.
Figure 2: Robust ASR pipeline, where enhanced speech $\hat{x}_{q}=\text{iSTFT}(\hat{X}_{q})$ is fed to backend ASR models for recognition. No joint training is performed. An optional speaker reinforcement module [72], which adds a scaled version of the input mixture signal $y_{q}$ to $\hat{x}_{q}$, can be included.
@@ 7 Experimental Setup
We do not use $\hat{X}_{q}$ to derive linear beamforming results for ASR [1, 33, 43], although this has been extremely popular, as we would like to validate whether the enhanced speech $\hat{X}_{q}$ itself is close to target speech and whether $\hat{X}_{q}$ itself can yield better ASR performance. We do not jointly train enhancement models with ASR models, as this requires knowledge of ASR models and would not accurately reflect the accuracy of $\hat{X}_{q}$ itself. We aim at building enhancement models that can produce enhanced speech with low distortion and high reduction to non-target signals. This way, the enhancement models could improve the robustness of many subsequent applications not limited to ASR.
In a nutshell, our main goal is to show, through SuperM2M, whether $\hat{X}_{q}$ itself would be better on real test data. We validate SuperM2M on CHiME-4 [42], a dataset consisting of simulated mixtures and real-recorded close-talk and far-field mixture pairs. To further show the effectiveness and potential of SuperM2M, a minor goal is to show whether SuperM2M can lead to state-of-the-art ASR performance on CHiME-4. The rest of this section describes the CHiME-4 dataset, miscellaneous system configurations, comparison systems, evaluation metrics, and several tricks to improve robust ASR performance.
@@ 7.1 CHiME-4 Dataset
CHiME-4 [42] is a major corpus for evaluating robust ASR and speech enhancement algorithms. It is recorded by using a tablet mounted with $6$ microphones, with the second microphone on the rear and the others facing front. The signals are recorded in four representative environments (including cafeteria, buses, pedestrian areas, and streets), where reverberation and directional, diffuse, transient and non-stationary noises naturally exist. During data collection, the target speaker hand-holds the tablet in a designated environment, and reads text prompts shown on the screen of the tablet. The target speaker wears a close-talk microphone so that the close-talk mixture can be recorded at the same time along with far-field mixtures recorded by the microphones on the tablet. The number of simulated and real-recorded utterances is listed in Table 1. It should be noted that, in CHiME-4, the room reverberation is weak, and the major challenge is in how to remove the multi-source non-stationary noise signals.
@@ 7.2 Evaluation Setup - Robust ASR
We check whether SuperM2M can improve ASR performance by feeding its enhanced speech to ASR models for decoding, following the pipeline in Fig. 2. We consider two ASR models.
Note that the WERs computed by ESPnet should not be directly compared with the ones by Whisper due to different text normalization. As a side result, to compare the ASR model in ESPnet with the Whisper model, we retrained an ESPnet model by using the same text normalization as the Whisper ASR model. We will report the evaluation results of using this ASR model later in Section 8.9.
@@ 7.5 Miscellaneous Configurations
TF-GridNet [18], which has shown strong separation performance in major benchmarks in supervised speech separation, is used as the DNN architecture. We consider two setups. Using the symbols defined in Table I of [18], the first one (denoted as TFGridNetv1) sets its hyper-parameters to $D=100$, $B=4$, $I=2$, $J=2$, $H=200$, $L=4$ and $E=2$, and the second one (denoted as TFGridNetv2) to $D=128$, $B=4$, $I=1$, $J=1$, $H=200$, $L=4$ and $E=4$. Please do not confuse these symbols with the ones in this paper. The models have $\sim$6.36.36.36.3 and $\sim$5.45.45.45.4 million parameters respectively. The v1 model uses approximately half of the computation and memory of v2888 This is because in v1, setting $I=2$ and $J=2$ cuts the sequence length to be modeled by the BLSTMs in TFGridNet by $50\%$, compared with setting $I=1$ and $J=1$ in v2. In this case, v1 would approximately use half of the computation and memory of v2 for the following reasons. First, in TFGridNet, the cross-frame self-attention module is much less costly than the sub-band temporal module and the intra-frame full-band module, both of which use BLSTMs, the most computationally-expensive block in each of the two modules. Second, although the input dimension to the BLSTMs is $200$ in v1 while $128$ in v2, the hidden dimensions of the BLSTMs in v1 and v2 are both set to $200$ and most computation inside BLSTMs is not spent between the input tensor and the hidden tensor. We highlight that this approximation is rough and is verified only empirically. , and is utilized for faster experimentation.
@@ 7.6 Comparison Systems
We consider the same DNN model trained only on the simulated data of CHiME-4 via supervised learning as the major baseline for comparison. We use exactly the same configurations as that in SuperM2M for traning. We denote this baseline as Supervised, to differentiate it with SuperM2M. Since CHiME-4 is a public and popular dataset, many existing models can be used directly for comparison.
@@ 7.7.1 SNR Augmentation for Simulated Training Mixtures
(#input mics: $6$; ASR model: Whisper Large v2)
@@
4e SuperM2M S+R TFGridNetv1 5 6+1 - 22.278 870 459 758 87422.27887045975887422.278\,870\,459\,758\,87422.278 870 459 758 874 22.600 104 856 813 6122.6001048568136122.600\,104\,856\,813\,6122.600 104 856 813 61 3.070 647 907 076 463.070647907076463.070\,647\,907\,076\,463.070 647 907 076 46 0.984 785 720 375 6820.9847857203756820.984\,785\,720\,375\,6820.984 785 720 375 682 3.517 263 633 430 138 43.51726363343013843.517\,263\,633\,430\,138\,43.517 263 633 430 138 4 3.965 148 642 652 575 33.96514864265257533.965\,148\,642\,652\,575\,33.965 148 642 652 575 3 4.052 936 311 000 8284.0529363110008284.052\,936\,311\,000\,8284.052 936 311 000 828 4.156 972 235 148 1314.1569722351481314.156\,972\,235\,148\,1314.156 972 235 148 131
4f SuperM2M S+R TFGridNetv1 6 6+1 - 22.339 721 592 086 73222.33972159208673222.339\,721\,592\,086\,73222.339 721 592 086 732 22.692 359 542 666 7922.6923595426667922.692\,359\,542\,666\,7922.692 359 542 666 79 3.101 081 421 429 4613.1010814214294613.101\,081\,421\,429\,4613.101 081 421 429 461 0.984 796 130 298 721 50.98479613029872150.984\,796\,130\,298\,721\,50.984 796 130 298 721 5 3.606 001 936 108 422 33.60600193610842233.606\,001\,936\,108\,422\,33.606 001 936 108 422 3 3.944 980 033 076 523.944980033076523.944\,980\,033\,076\,523.944 980 033 076 52 3.975 392 886 683 209 33.97539288668320933.975\,392\,886\,683\,209\,33.975 392 886 683 209 3 4.213 846 233 390 2084.2138462333902084.213\,846\,233\,390\,2084.213 846 233 390 208
4g SuperM2M S+R TFGridNetv1 7 6+1 - 22.205 550 543 286 58422.20555054328658422.205\,550\,543\,286\,58422.205 550 543 286 584 22.523 039 187 784 7922.5230391877847922.523\,039\,187\,784\,7922.523 039 187 784 79 3.096 138 221 567 327 43.09613822156732743.096\,138\,221\,567\,327\,43.096 138 221 567 327 4 0.984 329 144 290 849 90.98432914429084990.984\,329\,144\,290\,849\,90.984 329 144 290 849 9 3.529 364 311 068 086 63.52936431106808663.529\,364\,311\,068\,086\,63.529 364 311 068 086 6 3.993 384 696 059 053 83.99338469605905383.993\,384\,696\,059\,053\,83.993 384 696 059 053 8 4.006 410 256 410 2564.0064102564102564.006\,410\,256\,410\,2564.006 410 256 410 256 4.244 868 414 249 5224.2448684142495224.244\,868\,414\,249\,5224.244 868 414 249 522
4h SuperM2M S+R TFGridNetv1 8 6+1 - 22.356 131 915 793 27422.35613191579327422.356\,131\,915\,793\,27422.356 131 915 793 274 22.651 026 868 914 45722.65102686891445722.651\,026\,868\,914\,45722.651 026 868 914 457 3.135 788 249 608 242 33.13578824960824233.135\,788\,249\,608\,242\,33.135 788 249 608 242 3 0.984 945 777 711 785 30.98494577771178530.984\,945\,777\,711\,785\,30.984 945 777 711 785 3 3.509 196 515 004 843.509196515004843.509\,196\,515\,004\,843.509 196 515 004 84 3.928 845 145 415 6753.9288451454156753.928\,845\,145\,415\,6753.928 845 145 415 675 3.965 053 763 440 863.965053763440863.965\,053\,763\,440\,863.965 053 763 440 86 4.193 164 779 483 9984.1931647794839984.193\,164\,779\,483\,9984.193 164 779 483 998
@@
7a SuperM2M S+R TFGridNetv2 est. 6+1 - 22.548 556 464 188 02822.54855646418802822.548\,556\,464\,188\,02822.548 556 464 188 028 22.872 661 312 047 7722.8726613120477722.872\,661\,312\,047\,7722.872 661 312 047 77 3.152 997 804 500 839 83.15299780450083983.152\,997\,804\,500\,839\,83.152 997 804 500 839 8 0.985 174 543 509 160 90.98517454350916090.985\,174\,543\,509\,160\,90.985 174 543 509 160 9 3.505 162 955 792 1913.5051629557921913.505\,162\,955\,792\,1913.505 162 955 792 191 3.900 609 092 009 1973.9006090920091973.900\,609\,092\,009\,1973.900 609 092 009 197 3.866 832 092 638 5443.8668320926385443.866\,832\,092\,638\,5443.866 832 092 638 544 4.141 461 144 718 4744.1414611447184744.141\,461\,144\,718\,4744.141 461 144 718 474
7b SuperM2M S+R TFGridNetv2 est. 6+1 - 22.706 421 535 123 4822.7064215351234822.706\,421\,535\,123\,4822.706 421 535 123 48 23.159 114 919 114 66423.15911491911466423.159\,114\,919\,114\,66423.159 114 919 114 664 3.237 787 585 457 1663.2377875854571663.237\,787\,585\,457\,1663.237 787 585 457 166 0.986 333 159 671 4220.9863331596714220.986\,333\,159\,671\,4220.986 333 159 671 422 3.428 525 330 751 855 33.42852533075185533.428\,525\,330\,751\,855\,33.428 525 330 751 855 3 3.836 069 541 365 818 63.83606954136581863.836\,069\,541\,365\,818\,63.836 069 541 365 818 6 3.866 832 092 638 5443.8668320926385443.866\,832\,092\,638\,5443.866 832 092 638 544 3.981 179 876 945 3493.9811798769453493.981\,179\,876\,945\,3493.981 179 876 945 349
7c SuperM2M S+R TFGridNetv2 est. 6+1 - 22.849 200 136 282 22822.84920013628222822.849\,200\,136\,282\,22822.849 200 136 282 228 23.214 971 050 443 76523.21497105044376523.214\,971\,050\,443\,76523.214 971 050 443 765 3.215 787 200 223 3163.2157872002233163.215\,787\,200\,223\,3163.215 787 200 223 316 0.986 786 352 625 754 60.98678635262575460.986\,786\,352\,625\,754\,60.986 786 352 625 754 6 3.384 156 179 412 7143.3841561794127143.384\,156\,179\,412\,7143.384 156 179 412 714 3.807 833 487 959 340 53.80783348795934053.807\,833\,487\,959\,340\,53.807 833 487 959 340 5 3.773 779 983 457 4033.7737799834574033.773\,779\,983\,457\,4033.773 779 983 457 403 4.043 224 238 663 9784.0432242386639784.043\,224\,238\,663\,9784.043 224 238 663 978